# Uncertainty-Aware Adaptation of Large Language Models for Protein-Protein Interaction Analysis

### v1

@@ I Introduction
Traditional experimental methods for identifying PPIs, such as yeast two-hybrid screening and co-immunoprecipitation, have generated vast amounts of validated interaction data. These experimentally determined interactions have been systematically collected in comprehensive databases such as STRING [3], BioGRID [4], and IntAct [5], creating valuable resources for the research community. While experimental methods remain the gold standard for PPI validation, their labor-intensive and time-consuming nature has motivated the development of computational approaches to predict novel protein interactions.
Early computational methods for PPI prediction primarily relied on sequence-based evolutionary patterns across species [6, 7, 8]. As the field advanced, machine learning (ML) approaches emerged, offering new ways to integrate multiple types of biological data [9]. These included convolutional neural networks (CNNs) for analyzing protein sequence patterns, recurrent neural networks (RNNs) for capturing sequential dependencies, and graph neural networks (GNNs) for modeling the topology of protein interaction networks [10].
The advancements of large language models (LLMs) [11, 12, 13, 14, 15, 16, 17] have transformed multiple scientific domains through their unprecedented capabilities in understanding complex patterns and relationships in text data [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Building on their powerful language-processing capabilities, biology-specific models have been developed to tackle diverse tasks: ProtBERT [28] and ESM3 [29] focus on protein sequence analysis, while BioGPT [19] and BioMedGPT [30] have shown promise in extracting biological knowledge from scientific literature. Recently, [31] has investigated PPI prediction using LLM-based approaches, demonstrating the potential of these models in specialized biomedical tasks. However, a critical challenge remains: such models often produce overly confident predictions, especially when trained on limited data, posing serious risks in high-stakes biomedical applications [32, 33, 34, 35]. Uncertainty-aware adaptation of LLMs is particularly critical in biomedical applications where miscalibrated confidence levels could lead to erroneous conclusions about disease mechanisms or therapeutic targets. By incorporating principled uncertainty estimation and confidence calibration techniques, we can ensure that LLM-driven predictions are not only powerful but also reliable for guiding biomedical discoveries.
@@
Within the broader machine learning community, uncertainty quantification has long been a challenge due to the high-dimensional parameter spaces of deep learning models. Traditional Bayesian methods relying on Markov chain Monte Carlo sampling often become intractable at scale [36, 37]. Consequently, approximate Bayesian methods such as variational inference [38], sparse learning [39, 40], and dimension reduction [41] have been explored. Alternately, ensemble-based methods like deep ensembles [42], SWAG [43], and SeBayS [44] also provide principled strategies for capturing predictive variability. On the other hand, parameter-efficient fine-tuning techniques—particularly Low-Rank Adaptation (LoRA) [45]—makes it more tractable to adapt large models for specific downstream tasks. To this end, novel UQ-aware fine-tuning approaches have emerged that combine LoRA with Bayesian or ensemble-based ideas, such as Bayesian LoRA [46, 47, 48, 49] and LoRA ensembles [50, 51].
@@ II-B Bayesian model formulation
Let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1,\cdots,N}$ be a training dataset with $N$ i.i.d. samples, where $\bm{x}$ represents the input samples and $y$ represents the output samples. Bayesian inference captures model uncertainty by inferring a probability distribution over model parameters, $\bm{\theta}=(\theta_{1},\cdots,\theta_{T})\in\mathbb{R}^{T}$, instead of learning a single deterministic model–$p(y|\bm{x},\bm{\theta})$. The posterior distribution follows Bayes’ rule: $p(\bm{\theta}|\mathcal{D})\propto p(\mathcal{D}|\bm{\theta})p(\bm{\theta})$, where $p(\mathcal{D}|\bm{\theta})$ is the model likelihood and $p(\bm{\theta})$ is the prior distribution. To make predictions for a new input $\bm{x}_{\rm new}$, Bayesian model averaging (BMA) is applied using the posterior distribution $p(\bm{\theta}|\mathcal{D})$ as follows:
@@ III-A LoRA Ensemble
We employ an ensemble of LoRA models – LoRA Ensemble [50, 51] as an efficient strategy for uncertainty quantification in LLMs. Traditional deep ensembles yield better predictive performance and uncertainty estimation by training multiple models independently, but applying this directly to LLMs is often infeasible due to high memory and computational costs.
To circumvent these issues, each LoRA Ensemble member fine-tunes the same pre-trained backbone $\bm{W}_{0}$ with a low-rank trainable modification $\Delta\bm{W}_{m}=\bm{B}_{m}\bm{A}_{m}$, where $\bm{B}_{m}\in\mathbb{R}^{d_{1}\times r}$ and $\bm{A}_{m}\in\mathbb{R}^{r\times d_{2}}$ have significantly fewer parameters than the full model, $r_{m}\ll\min(d_{1},d_{2})$. These adapters are trained independently and in parallel, ensuring diverse solutions–$\{\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{M}\}$. The ensemble prediction is computed by averaging outputs across $M$ ensemble members. For a given input $\bm{x}_{\rm new}$, if $y^{m}_{\rm new}$ represents the prediction from the $m$-th ensemble member, the final ensemble output (for continuous outcomes) is given by:
@@ III-B Bayesian Low-Rank Adaptation
$=\underset{\bm{\theta}}{\arg\!\max}\log{\rm P}(\bm{y}|\bm{X},\bm{\theta})+\frac{\lambda}{2}||\bm{\theta}||^{2}_{2}+{\rm const}$
where $\bm{X}$ represents the model inputs. The term associated with log of the prior distribution provides us $L_{2}$-regularization on the trainable parameters. We can incorporate this in frequentist model training via weight decay term with $\lambda/2$ strength. As a result, parameters from any previously trained model that used a reasonable weight decay setting (for example, via AdamW with its weight decay) can be directly reused.
@@
$\bm{F}(\bm{\theta})=\sum_{n}^{N}\mathbb{E}_{\hat{y}\sim{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}\bm{G}^{\top}\bigr{]}$
@@
$\bm{F}_{l}=\sum_{n=1}^{N}\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{a}_{l-1}\bm{a}_{l-1}^{\top}\bigr{]}\otimes\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}_{\bm{s}_{l}}\bm{G}_{\bm{s}_{l}}^{\top}\bigr{]}$ (2)
This approach avoids storing the full, dense Hessian, thereby reducing computational overhead. By applying K-FAC to the LoRA parameters, we maintain a compact representation of uncertainty while keeping the overhead similar to standard training. However, in Equation (2), the first expectation grows with the square of the layer’s input width, while the second grows with the square of the output width. Because LoRA adapters alternate between wide-input-narrow-output configuration and vice versa, one of these expectations can become especially large. To address this, we use an incremental SVD to factorize the large matrix into two new low-rank factors thereby saving memory. Further mathematical details are provided in Appendix E of [46].
@@
$f_{\bm{\theta}}(\bm{x}_{\rm new})\approx f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})+\nabla_{\bm{\theta}}f_{\bm{\theta}}(\bm{x}_{\rm new})\bigl{|}_{\bm{\theta}_{\rm MAP}}^{\top}\bigl{(}\bm{\theta}-\bm{\theta}_{\rm MAP}\bigr{)}.$
@@
wh ere $\text{ere }\Lambda=\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})^{\top}H^{-1}\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new}).$
@@
TABLE I: ND-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 86.510.54 0.3620.036 0.0950.011 0.9660.005 0.8810.005 0.8630.006 0.7450.011 0.9530.004
LoRA Ensemble 88.700.62 0.3020.002 0.0880.005 0.9730.001 0.8990.004 0.8860.006 0.7850.011 0.9640.001
Bayesian LoRA 86.510.20 0.3170.003 0.0520.021 0.8480.033 0.8660.002 0.8650.002 0.7320.003 0.9440.004
BioMedGPT Single LoRA 85.442.16 0.5390.053 0.1190.025 0.9630.016 0.8740.011 0.8520.023 0.7260.034 0.9440.005
LoRA Ensemble 88.001.19 0.3630.049 0.0870.022 0.9650.008 0.8920.007 0.8790.013 0.7710.019 0.9560.001
Bayesian LoRA 86.820.60 0.3200.012 0.0330.007 0.8690.015 0.8680.006 0.8680.006 0.7370.012 0.9370.003
TABLE II: M-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 85.820.26 0.3980.016 0.0840.006 0.9080.036 0.8630.007 0.8580.002 0.7210.009 0.9370.002
@@
BioMedGPT Single LoRA 83.680.54 0.5420.026 0.1130.009 0.7990.047 0.8400.002 0.8370.006 0.6780.007 0.9250.005
LoRA Ensemble 87.141.39 0.3540.028 0.0620.010 0.8640.036 0.8720.013 0.8710.014 0.7440.027 0.9410.005
Bayesian LoRA 83.290.57 0.3850.015 0.0370.018 0.8880.033 0.8380.003 0.8320.007 0.6710.007 0.9050.004
TABLE III: C-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 96.620.62 0.0940.011 0.0330.002 0.9730.016 0.9670.007 0.9660.006 0.9320.013 0.9960.002
@@
BioMedGPT Single LoRA 97.680.82 0.0590.011 0.0250.011 0.9760.006 0.9770.008 0.9770.008 0.9540.016 0.9980.001
@@ IV Experimental Results
In this section, we assess the performance of two uncertainty-aware LoRA adaptations—LoRA Ensemble and Bayesian LoRA—applied to LLaMA-3-8B and BioMedGPT-LM-7B models on publicly available protein-protein interaction datasets. As a baseline, we include a single LoRA model trained in a deterministic manner. All LoRA-based approaches were implemented using the PEFT library [54], with each configuration run three times using different random seeds. We evaluate model performance and robustness by accuracy (Acc), negative log-likelihood (NLL), and expected calibration error (ECE) on the test sets. Additional details on the NLL and ECE metrics can be found in Appendix V-A. Furthermore, we report Matthews Correlation Coefficient (MCC), specificity (Spec.), precision (Prec.), F1-score, and Area under Receiver Operating Characteristic curve (AUROC) over test sets for a comprehensive view of predictive capabilities. Final metrics are summarized by the mean and standard deviation across three independent runs.
PPI Datasets. The datasets analyzed here explore PPIs related to various diseases, providing valuable insights into their underlying mechanisms. The Neurodegenerative diseases PPI (ND-PPI) dataset, sourced from the study [55], focuses on neurodegenerative diseases and examines a network of 820 proteins forming 11,762 interactions, evenly split between positive and negative pairs. The dataset is structured to assess whether specific protein pairs interact in the presence of neurodegenerative conditions. Similarly, the metabolic disorders PPI (M-PPI) dataset, also from [55], investigates metabolic disorders, encompassing 1,063 proteins and a total of 10,262 interactions. The cancer PPI (C-PPI) dataset, derived from the study [56], consisted of 933 positive and 1,308 negative interactions. To ensure balanced representation, this dataset was curated to create an equal-sized collection of 1,866 total interactions. These datasets have been evaluated by prompting models to assess whether proteins interact in the corresponding conditions and collectively contribute to advancing computational models for predicting PPIs across different disease contexts, enhancing our understanding of disease-specific interaction networks. These tasks are formalized into binary (True/False) classification problems as illustrated in Fig. 1. Furthermore, each dataset is divided into $80\%$ for training and $20\%$ for testing, with all models evaluated on the fixed test set in each PPI prediction task. We refer readers to [31], for additional details and exploratory analyses of these datasets.
@@
Implementation Details. In all experiments, we construct a LoRA ensemble using three individually fine-tuned LoRA learners. The LoRA matrices $\bm{B}$ are initialized to zero, while the entries of $\bm{A}$ follow a Kaiming Uniform initialization [57]. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, default hyperparameters, and a total of four training epochs. The batch size is set to $4$ for the ND-PPI and M-PPI cases and $16$ for the C-PPI case, following [31]. For Bayesian LoRA, the prior precision $\lambda$ is fixed at $0.1$. Lastly, LoRA is applied to the queries, values, and output layer across all methods, with specific hyperparameters set to $r=16$, $\alpha=32$, a dropout rate of $0.05$, and a maximum sequence length of $50$.
Results. The results for ND-PPI, M-PPI, and C-PPI tasks are summarized in Tables I, II, and III, respectively.
@@
In the ND-PPI prediction task (Table I), we demonstrate that the LoRA ensemble achieves the highest predictive accuracy among all models in both LLM settings and has the lowest NLL in the LLaMA-3 fine-tuning case. Conversely, Bayesian LoRA demonstrates the best calibration in both scenarios, exhibiting the lowest ECE and achieving the lowest NLL in the BioMedGPT fine-tuning case. Lastly, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. In the M-PPI prediction task (Table II), we show that the LoRA ensemble achieves the highest predictive accuracy and lowest NLL in both LLM scenarios, while also attaining the lowest ECE in the LLaMA-3 case. Conversely, Bayesian LoRA achieves the best calibration in the BioMedGPT case and the highest specificity in both scenarios. Finally, the LoRA ensemble outperforms all the models by achieving best precision, F1-score, MCC, and AUROC values.
@@
In the C-PPI prediction task (Table III), we demonstrate that the LoRA ensemble once again achieves the highest predictive accuracy and lowest NLL in both settings, while also attaining the lowest ECE in the BioMedGPT scenario. Bayesian LoRA matches the best predictive accuracy in the BioMedGPT case and achieves the lowest ECE in the LLaMA-3 case. In the LLaMA-3 setting, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. Additionally, it achieves the best specificity in the BioMedGPT case. Notably, both Bayesian LoRA and the LoRA ensemble attain the best precision, F1-score, and MCC values in the BioMedGPT case. Lastly, all three models yield identical AUROC values in the BioMedGPT case.

### v2

@@ I Introduction
Traditional experimental methods for identifying PPIs, such as yeast two-hybrid screening and co-immunoprecipitation, have generated vast amounts of validated interaction data. These experimentally determined interactions have been systematically collected in comprehensive databases such as STRING [3], BioGRID [4], and IntAct [5], creating valuable resources for the research community. While experimental methods remain the gold standard for PPI validation, their labor- and time-intensive nature has led to the development of computational approaches to predict protein interactions.
Early computational methods for PPI prediction primarily relied on sequence-based evolutionary patterns across species [6, 7, 8]. As the field advanced, machine learning (ML) approaches emerged, offering new ways to integrate multiple types of biological data [9, 10]. These included convolutional neural networks for analyzing protein sequence patterns, recurrent neural networks for capturing sequential dependencies, and graph neural networks for modeling the topology of protein interaction networks [11].
The advancements of large language models (LLMs) [12, 13, 14, 15, 16, 17] have transformed multiple scientific domains through their unprecedented capabilities in understanding complex patterns and relationships in text data [18, 19, 20, 21, 22, 23, 24, 25]. Building on their powerful language-processing capabilities, biology-specific models have been developed to tackle diverse tasks: ProtBERT [26] and ESM3 [27] focus on protein sequence analysis, while BioGPT [19] and BioMedGPT [28] have shown promise in extracting biological knowledge from scientific literature. Recently, [29] has investigated PPI prediction using LLM-based approaches, demonstrating the potential of these models in specialized biomedical tasks. However, a critical challenge remains: such models often produce overly confident predictions, especially when trained on limited data, posing serious risks in high-stakes biomedical applications [30, 31, 32, 33]. Uncertainty-aware adaptation of LLMs is particularly critical in biomedical applications where miscalibrated confidence levels could lead to erroneous conclusions about disease mechanisms or therapeutic targets. By incorporating principled uncertainty estimation and confidence calibration techniques, we can ensure that LLM-driven predictions are not only powerful but also reliable for guiding biomedical discoveries.
Within the broader machine learning community, uncertainty quantification has long been a challenge due to the high-dimensional parameter spaces of deep learning models. Traditional Bayesian methods relying on Markov chain Monte Carlo sampling often become intractable at scale [34, 35]. Consequently, approximate Bayesian methods such as variational inference [36], sparse learning [37, 38], and dimension reduction [39] have been explored. Alternately, ensemble-based methods like deep ensembles [40], SWAG [41], and SeBayS [42] also provide principled strategies for capturing predictive variability. On the other hand, parameter-efficient fine-tuning techniques—particularly Low-Rank Adaptation (LoRA) [43]—makes it tractable to adapt large models for specific downstream tasks. To this end, novel UQ-aware fine-tuning approaches have emerged that combine LoRA with Bayesian or ensemble ideas, such as Bayesian LoRA [44, 45, 46, 47] and LoRA ensembles [48, 49].
@@ III-A LoRA Ensemble
We employ an ensemble of LoRA models – LoRA Ensemble [48, 49] as an efficient strategy for uncertainty quantification in LLMs. Traditional deep ensembles yield better predictive performance and uncertainty estimation by training multiple models independently, but applying this directly to LLMs is infeasible due to high memory and computational costs.
[…]
@@ III-B Bayesian Low-Rank Adaptation
[…]
where $\bm{X}$ represents the model inputs. The log prior term yields $L_{2}$-regularization on the trainable parameters, which can be incorporated into frequentist model training as a weight decay term with strength $\lambda/2$. Consequently, parameters from any model trained with an appropriate weight decay setting (e.g., via AdamW with its weight deca) can be directly reused.
@@
$\bm{F}(\bm{\theta})=\sum_{n}^{N}\mathbb{E}_{\hat{y}\sim{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}\bm{G}^{\top}\bigr{]}\vskip-1.99997pt$
@@
$\bm{F}_{l}=\sum_{n=1}^{N}\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{a}_{l-1}\bm{a}_{l-1}^{\top}\bigr{]}\otimes\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}_{\bm{s}_{l}}\bm{G}_{\bm{s}_{l}}^{\top}\bigr{]}\vskip-1.99997pt$ (2)
This approach avoids storing the full, dense Hessian, thereby reducing computational overhead. By applying K-FAC to the LoRA parameters, we maintain a compact representation of uncertainty while keeping the overhead similar to standard training. However, in Equation (2), the first expectation grows with the square of the layer’s input width, while the second grows with the square of the output width. Because LoRA adapters alternate between wide-input-narrow-output configuration and vice versa, one of these expectations can become especially large. To address this, we use an incremental singular value decomposition to factorize the large matrix into two new low-rank factors thereby saving memory. Further mathematical details are provided in Appendix E of [44].
@@
$f_{\bm{\theta}}(\bm{x}_{\rm new})\approx f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})+\nabla_{\bm{\theta}}f_{\bm{\theta}}(\bm{x}_{\rm new})\bigl{|}_{\bm{\theta}_{\rm MAP}}^{\top}\bigl{(}\bm{\theta}-\bm{\theta}_{\rm MAP}\bigr{)}.\vskip-1.99997pt$
@@
wh $\text{ere }\Lambda=\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})^{\top}H^{-1}\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new}).$
@@
TABLE I: PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Dataset Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
LLaMA-3 ND-PPI Single LoRA 87.421.64 0.3690.040 0.0890.023 0.9360.049 0.8850.007 0.8730.017 0.7580.025 0.9560.001
LoRA Ensemble 89.840.30 0.2690.007 0.0540.002 0.9590.005 0.9050.002 0.8980.003 0.8020.005 0.9630.001
Bayesian LoRA 86.870.73 0.3140.005 0.0310.005 0.8720.031 0.8690.008 0.8690.007 0.7380.015 0.9400.002
M-PPI Single LoRA 85.820.26 0.3980.016 0.0840.006 0.9080.036 0.8630.007 0.8580.002 0.7210.009 0.9370.002
@@
C-PPI Single LoRA 96.620.62 0.0940.011 0.0330.002 0.9730.016 0.9670.007 0.9660.006 0.9320.013 0.9960.002
@@
BioMedGPT ND-PPI Single LoRA 86.900.99 0.5270.093 0.1010.013 0.9480.007 0.8790.008 0.8680.010 0.7470.018 0.9440.004
LoRA Ensemble 88.940.75 0.3660.021 0.0730.011 0.9560.005 0.8970.006 0.8890.008 0.7850.013 0.9560.001
Bayesian LoRA 87.170.36 0.3150.003 0.0310.003 0.8820.019 0.8720.004 0.8720.004 0.7440.008 0.9380.001
M-PPI Single LoRA 85.990.44 0.4780.008 0.0920.003 0.8950.037 0.8630.007 0.8600.004 0.7220.011 0.9340.001
LoRA Ensemble 87.451.20 0.3470.006 0.0530.001 0.9150.013 0.8770.011 0.8740.012 0.7510.023 0.9430.002
Bayesian LoRA 83.370.74 0.3820.014 0.0330.011 0.8680.019 0.8360.006 0.8330.008 0.6690.013 0.9070.005
C-PPI Single LoRA 97.680.82 0.0590.011 0.0250.011 0.9760.006 0.9770.008 0.9770.008 0.9540.016 0.9980.001
@@ IV Experimental Results
In this section, we assess the performance of two uncertainty-aware LoRA adaptations—LoRA Ensemble and Bayesian LoRA—applied to LLaMA-3-8B and BioMedGPT-LM-7B models on publicly available protein-protein interaction datasets. As a baseline, we include a single LoRA model trained in a deterministic manner. All LoRA-based approaches were implemented using the PEFT library [52], with each configuration run three times using different random seeds. We evaluate model performance and robustness by accuracy (Acc), negative log-likelihood (NLL), and expected calibration error (ECE) on the test sets. Additional details on the NLL and ECE metrics can be found in Appendix-A. Furthermore, we report Matthews Correlation Coefficient (MCC), specificity (Spec.), precision (Prec.), F1-score, and Area under Receiver Operating Characteristic curve (AUROC) over test sets for a comprehensive view of predictive capabilities. Final metrics are summarized by the mean and standard deviation across three independent runs.
PPI Datasets. The datasets analyzed in this work focus on PPIs relevant to various disease contexts, offering important insights into the underlying molecular mechanisms. We employ three primary balanced datasets—Neurodegenerative diseases PPI (ND-PPI), Metabolic disorders PPI (M-PPI), and Cancer PPI (C-PPI)-all of them containing equal numbers of positive and negative interactions.
ND-PPI Dataset. Adapted from [53] and [29], this dataset comprises 820 proteins involved in neurodegenerative diseases (e.g., Alzheimer’s, Parkinson’s, Huntington’s). Positive interactions are sourced from experimental and literature data, while negative pairs are generated by selecting protein pairs with no documented evidence of interaction based on established databases (e.g., DIP, BioGRID) with similar characteristics (e.g., similar functional annotations or expression profiles), resulting in 11,762 balanced interactions.
M-PPI Dataset. Also detailed in [53] and [29], the M-PPI dataset covers 1,063 proteins related to metabolic disorders (e.g., diabetes, obesity). As with the ND-PPI dataset, experimentally supported PPIs formed the positive set, while negative pairs were systematically identified based on the absence of known interactions. Consistent sampling methods ensure that protein abundance and disease relevance are preserved, resulting in a balanced set of 10,262 interactions.
C-PPI Dataset. Derived from [54], the cancer PPI dataset originally had 933 positive and 1,308 negative interactions. Negative pairs were curated by excluding proteins with known interactions, and the set was balanced to 1,866 total interactions, in line with symmetric logistic matrix factorization approach [53].
These datasets are employed to evaluate and advance computational models for predicting protein–protein interactions across distinct disease contexts by framing the tasks as binary (True/False) classification problems (see Fig. 1). Models assess whether proteins interact under specific conditions, contributing to our understanding of disease-specific interaction networks. Each dataset is partitioned into an 80% training set and a 20% testing set, consistent with network-based PPI prediction frameworks [54] and recent advances in large language model analysis under diverse biological conditions [29]. For further details on dataset selection, exploratory analyses, and additional context-specific investigations, readers are referred to [29] and related work.
@@
Implementation Details. In all experiments, we construct a LoRA ensemble using three individually fine-tuned LoRA learners. The LoRA matrices $\bm{B}$ are initialized to zero, while the entries of $\bm{A}$ follow a Kaiming Uniform initialization [55]. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, default hyperparameters, and a total of four training epochs. The batch size is set to $4$ for the ND-PPI and M-PPI cases and $16$ for the C-PPI case, following [29]. For Bayesian LoRA, the prior precision $\lambda$ is fixed at $0.1$. We select the LoRA rank $(r)$ via hyperparameter search (Appendix-B) and report results for each LLM backbone and PPI dataset combination in Table I using the finalized $r$. Finally, LoRA adaptations are applied to the query, value, and output layer across all methods with $\alpha=32$, a dropout rate of $0.05$, and a maximum sequence length of $50$.
Results. Table I summarizes the performance of the three LoRA-based models on ND-PPI, M-PPI, and C-PPI tasks using LLaMA-3 and BioMedGPT backbones.
@@
In the ND-PPI prediction task, we demonstrate that the LoRA ensemble achieves the highest predictive accuracy among all models in both LLM settings and has the lowest NLL in the LLaMA-3 fine-tuning case. Conversely, Bayesian LoRA demonstrates the best calibration in both scenarios, exhibiting the lowest ECE and achieving the lowest NLL in the BioMedGPT fine-tuning case. Lastly, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. In the M-PPI prediction task, we show that the LoRA ensemble achieves the highest predictive accuracy and lowest NLL in both LLM scenarios, while also attaining the lowest ECE in the LLaMA-3 case and highest specificity in BioMedGPT case. Conversely, Bayesian LoRA achieves the best calibration in the BioMedGPT case and the highest specificity in LLaMA-3 case. Finally, the LoRA ensemble outperforms all the models by achieving best precision, F1-score, MCC, and AUROC values.
@@
In the C-PPI prediction task, we demonstrate that the LoRA ensemble once again achieves the highest predictive accuracy and lowest NLL in both settings, while also attaining the lowest ECE in the BioMedGPT scenario. Bayesian LoRA matches the best predictive accuracy in the BioMedGPT case and achieves the lowest ECE in the LLaMA-3 case. In the LLaMA-3 setting, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. Additionally, it achieves the best specificity in the BioMedGPT case. Notably, both Bayesian LoRA and the LoRA ensemble attain the best precision, F1-score, and MCC values in the BioMedGPT case. Lastly, all three models yield identical AUROC values in the BioMedGPT case.
We run one‐sided Welch’s t-tests to see if the best-performing uncertainty-aware LoRA variant (ensemble or Bayesian) outperforms Single LoRA in accuracy or lowers NLL/ECE. In most LLaMA-3 and BioMedGPT experiments—especially ND-PPI and M-PPI—the p-values are below 0.05, confirming significant gains; a handful of cases (LLaMA-3 ND-PPI accuracy, BioMedGPT M-PPI accuracy, and all BioMedGPT C-PPI metrics) are not significant, indicating comparable performance.
Next, in Fig. 2 we show reliability diagrams for LLaMA-3 on ND-PPI (top row), M-PPI (middle row), and C-PPI (bottom row), comparing Single LoRA (left column), LoRA Ensemble (middle column), and Bayesian LoRA (right column). These diagrams plot binned confidence versus observed accuracy with the perfect calibration following the diagonal identity line, and any highlighted gap indicates miscalibration. Both LoRA ensemble and Bayesian LoRA variants consistently reduce this gap (and achieve lower ECE) relative to the single LoRA. We have omitted BioMedGPT plots for brevity.
Figure 2: Reliability Diagrams: A visual representation of model calibration for the LLaMA-3 backbone trained on the ND-PPI (top row), M-PPI (middle row), and C-PPI (bottom row). Columns correspond, left to right, to the single LoRA, LoRA ensemble, and Bayesian LoRA models.
TABLE II: Effect of LoRA rank $(r)$ in single LoRA, LoRA ensemble, and Bayesian LORA models applied to Llama-3 and BioMedGPT across three PPI datasets. Final $r$ choices for each LLM backbone and dataset combination are shown by marking the corresponding results with light green color. Metrics are reported as means with standard deviations in subscript over three runs.
LLM Model Dataset Methods Rank=8=8 Rank=16=16 Rank=32=32
Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$)
LLaMA-3 ND-PPI Single LoRA 88.470.60 0.3300.036 0.0690.015 86.510.54 0.3620.036 0.0950.011 87.421.64 0.3690.040 0.0890.023
LoRA Ensemble 89.571.15 0.2780.022 0.0630.014 88.700.62 0.3020.002 0.0880.005 89.840.30 0.2690.007 0.0540.002
Bayesian LoRA 85.141.45 0.3300.020 0.0600.017 86.510.20 0.3170.003 0.0520.021 86.870.73 0.3140.005 0.0310.005
M-PPI Single LoRA 85.840.35 0.4340.040 0.0890.012 85.820.26 0.3980.016 0.0840.006 86.151.09 0.5150.072 0.1010.014
LoRA Ensemble 87.160.67 0.3340.012 0.0560.004 87.450.16 0.3080.013 0.0510.010 87.500.75 0.3380.018 0.0550.005
Bayesian LoRA 82.371.97 0.3920.038 0.0770.044 83.411.17 0.3740.005 0.0710.018 81.782.06 0.4050.051 0.0930.046
C-PPI Single LoRA 96.611.34 0.1160.024 0.0370.012 96.620.62 0.0940.011 0.0330.002 96.610.31 0.1450.057 0.0340.006
LoRA Ensemble 97.330.54 0.0900.020 0.0360.005 97.860.00 0.0660.005 0.0290.005 96.790.54 0.0810.019 0.0300.005
Bayesian LoRA 96.610.82 0.1420.025 0.0340.008 96.971.24 0.0850.020 0.0270.002 96.791.42 0.1190.040 0.0330.014
BioMedGPT ND-PPI Single LoRA 86.291.58 0.4710.095 0.0950.022 85.442.16 0.5390.053 0.1190.025 86.900.99 0.5270.093 0.1010.013
LoRA Ensemble 87.480.63 0.4030.065 0.0980.019 88.001.19 0.3630.049 0.0870.022 88.940.75 0.3660.021 0.0730.011
Bayesian LoRA 86.540.36 0.3260.007 0.0310.010 86.820.60 0.3200.012 0.0330.007 87.170.36 0.3150.003 0.0310.003
M-PPI Single LoRA 85.990.44 0.4780.008 0.0920.003 83.680.54 0.5420.026 0.1130.009 84.950.43 0.5100.060 0.1010.008
LoRA Ensemble 87.451.20 0.3470.006 0.0530.001 87.141.39 0.3540.028 0.0620.010 86.650.78 0.3570.003 0.0610.000
Bayesian LoRA 83.370.74 0.3820.014 0.0330.011 83.290.57 0.3850.015 0.0370.018 83.440.50 0.3740.005 0.0280.010
C-PPI Single LoRA 97.860.54 0.0790.015 0.0260.002 97.680.82 0.0590.011 0.0250.011 97.510.31 0.0630.005 0.0260.003
LoRA Ensemble 98.570.62 0.0630.009 0.0240.004 98.400.54 0.0520.000 0.0210.002 98.400.00 0.0570.002 0.0240.001
Bayesian LoRA 98.220.31 0.0720.009 0.0280.002 98.400.54 0.0640.005 0.0310.001 97.860.54 0.0750.010 0.0310.007